A baseball player gets 16 hits in 50 at bats. What percent are hits and what percent are nonhits.
Answer by mangopeeler07(462)

50 at bats
16 hits

50-16 nonhits
34 non hits

16/50=?
34/50=??

16/50=.32
34/50=.68

move the decimal twice to the right

hits=32%
non-hits=68%